# atomic orbital API Application Programmer Interface

### 06 Mar 2024 v3

@@
RKB
restricted kinetic balance
@@
SAPT
symmetry-adapted perturbation theory
SERS
surface-enhanced raman scattering
@@
ae
almost everywhere
BVP
boundary value problem
PDE
partial differential equation
RDM
1-body reduced density matrix
SCRF
self-consistent reaction field
IEFPCM
Integral Equation Formalism polarizable continuum model (PCM)
FMM
fast multipole method
DD
domain decomposition
TRS
time-reversal symmetry
SI
Supporting Information
DHF
Dirac–Hartree–Fock
MO
molecular orbital
[1,2]Christian Tantardini [1]Luca Frediani
1]Hylleraas center, Department of Chemistry, UiT The Arctic University of Norway, PO Box 6050 Langnes, N-9037 Tromsø, Norway. 2]Department of Materials Science and NanoEngineering, Rice University, Houston, Texas 77005, United States of America. 3]Algorithmiq Ltd., Kanavakatu 3C, FI-00160, Helsinki, Finland.
@@
christiantantardini@ymail.com    Roberto Di Remigio Eikås    luca.frediani@uit.no [ [ [
@@ Abstract
The many-body wave function of an $N$-electron system within a relativistic framework can be described by the Dirac equation. Unfortunately, the Dirac operator $\hat{\mathfrak{D}}$ is unbounded. Thus, anions show the variational collapse of $N+1$ electron wavefunction making impossible to extend the $\hat{\mathfrak{D}}$ operator to such systems. Thus, it is necessary to avoid it and an alternative approach is based on applying the square of the Dirac operator $\hat{\mathfrak{D}}^{2}$. This approach is especially suitable for a multiwavelet framework: its implementation in an integral equation form is readily available and the completeness of the multiwavelet basis affords tight error control. The use of the $\hat{\mathfrak{D}}^{2}$ operator within a multiwavelet (MW) framework is able to achieve higher precision compared to the $\hat{\mathfrak{D}}$, consistent with the quadratic nature of the error estimates in a variational framework.
@@
, multiwavelets, Dirac equation, variational collapse, squared-Dirac operator
@@ 1 Introduction
Dirac’s equation describes the relativistic motion of electrons and still today represents the most accurate physical model, short of quantum electrodynamics, for the inclusion of relativistic effects in molecular calculations. Paul Dirac formulated his equation by aiming for a quantum mechanical equation that satisfied Lorentz covariance, while guaranteeing a positive semi-definite probability density.[1]
In its original formulation, the Dirac equation suffers from some drawbacks and the most important one is called variational collapse, a term coined by Schwarz and Wallmeier [2], because the Dirac operator $\hat{\mathfrak{D}}$, in contrast to the Schrödinger Hamiltonian $\hat{H}$, is not bounded from below giving spurious unphysical solutions as the kinetic energy tends to zero. Therefore, the application of variational methods is not justified. Schwarz and Wallmeier [2] were the first to observe: (i) the collapsing of two neighboring exponents of basis sets to the same value during optimization; (ii) no systematic variation of the eigenvalues with increasing basis size (n.b. eigenvalues change with basis set). During the years the same was observed by others [3, 4, 5], and recently with contracted Dyall basis sets [6, 7, 8, 9, 10] and atomic-balanced basis sets [11, 12, 13].
Such effects were attenuated by expanding the basis sets through a Legendre parameterization which provides an efficient way of systematically approaching the basis set limit and by using the restricted kinetic balance (RKB) to generate the small components of the spinorbitals [14, 15, 16]. However, the presence of uncontracted basis sets with the RKB could generate the prolapse due to the lack of kinetic balance in the description of the negative-energy states [17, 14, 11, 18, 19].
Nowadays, the variational collapse can be considered largely mitigated by the use of appropriate atomic basis sets in conjunction with the RKB prescription, but the problem continues to exist for electronic metastable anions, where the four-component Dirac wavefunction of $N+1$ electron will collapse in an orbital that is infinitely removed from the parent plus forming a continuum function [20].
There are essentially four strategies to deal with the variational collapse:
  1. 1.
    accepting variational collapse and trying to minimize its effects by suppressing the small components and introducing $c\rightarrow\infty$ [3, 21, 22] or special choice of basis set [23];
  2. 2.
    transforming the Dirac operator (e.g., Foldy-Wouthuysen (FW) transformation [24, 25]) to achieve a matrix representation that incorporates the correct non-relativistic limit;
  3. 3.
    manipulating the matrix representation of the Dirac operator (e.g., by modifying $\vec{\sigma}\cdot\vec{p}$ [5], by the stabilization method [26, 27] or the rotation method [28]) without transforming the operator itself;
  4. 4.
    replacing the Dirac operator by a related operator which has the same eigenfunctions, but which is bounded from below such that the variation principle can be applied (e.g., squared Dirac operator [4] or projection of the Dirac operator to positive energy states [29, 30]).
Here, we will pursue the latter option in the field of adaptive multiwavelet basis. The use of $\hat{\mathfrak{D}}^{2}$ was originally presented by Wallmeier and Kutzelnigg [4] for one-electron calculations and later extended to molecular calculations [31]. We first review the theory of $\hat{\mathfrak{D}}^{2}$ for one- and many-electron systems. For the latter, we introduce the unretarded Coulombic interaction to describe the repulsion between electrons, leading to the direct ($\hat{J}$) and exchange ($\hat{K}$) interactions in the mean-field case. We then apply this method to simple one- and two-electron systems to show the viability of the approach and the precision which can be attained compared to the standard Dirac equation.
@@ 2 Theoretical Background
We use indices $u,w\in\{x,y,z\}$ for Cartesian components. The $2\times 2$ matrices $\sigma_{u},\,u\in\{x,y,z\}$ appearing as anti-diagonal blocks are the Pauli matrices. The $4\times 4$ matrix $\beta$ is defined as:
@@
The large and small components can be transformed into each other. Practically, the exact relationship is only applicable for the free-particle problem. In atomic and molecular applications, one initializes the small components by applying the kinetic balance condition, i.e. the nonrelativistic limit of the exact large-small relationship, to the large components [32]. In the following, the four-component molecular orbitals will further be indexed with a subscript Latin letters, e.g. $p,q,i,j,\ldots\in\{1,\ldots,N\}$, with $N$ the number of such orbitals.
@@ 2.1 Adaptive Multiwavelets Basis
MRA is a numerical framework devised to solve integral and differential equations with arbitrary, predefined precision.[33] Standard quantum chemical methods expand the unknown molecular orbitals into fixed atom-centered bases and recast the solution of the Hartree–Fock equations as a matrix eigenvalue problem to determine the expansion coefficients [34]. multiresolution analysis (MRA)-based methods instead solve directly for the occupied MOs, through adaptive refinement of the expansion basis. The mathematical framework enabling such adaptive algorithms relies on the construction of a sequence of orthogonal subspaces which is dense in the space of square-integrable functions for a user-defined simulation box.
@@
where $l=0,\ldots,2^{n-1}$ is the translation index, identifying functions in the sub-interval $[l/2^{n},(l+1)/2^{n}]$. This construction defines a telescopic series of scaling spaces:
@@
$S_{n}^{k}\oplus W_{n}^{k}=S_{n+1}^{k},\>\>\>W_{n}^{k}\perp S_{n}^{k},$ (9)
The crucial property of wavelet functions is constituted by their vanishing moments, which are a straightforward consequence of their construction as orthogonal complement of a set of polynomial functions. This is the key feature that leads to sparse representations of functions, enabling both precision control on the one hand and fast algorithms on the other. The $k$ parameter is the order of the Legendre polynomials used in each refinement box, which is regardless of the number of of electrons in the system and, in principle, a given a user-defined tolerance, $\epsilon$, which is used both as requested precision in the adaptive refinement and operator application algorithms, and as convergence tolerance of the self-consistent field (SCF) cycle. Convergence is achieved once the norm of the orbital update drops below $\epsilon$. This means that if the error on the orbital is $10^{-N}$ the error on the energy will be $10^{-2N}$.
We note that the MRA construction is dense in $L^{2}$ not because of the high polynomial order, but because of the telescopic construction of subspaces by dyadic subdivision. In principle, one could achieve completeness of the representation basis even using $k=1$, the telescopic series of spaces would however be impractically large. In practice, as we showed in our previous work on the full Breit Hamiltonian [35], requesting tighter precision in a relativistic setting demands a higher polynomial order than what provided by default by MRCPP, our MRA library. The interested reader is referred to the available literature for details about MRA and its practical realization for quantum chemistry applications.[36, 37, 38, 39, 40]
@@ 2.2 Integral equation formulation for the Dirac operator
where $\hat{h}_{D}=c\vec{\alpha}\cdot\vec{p}+\beta mc^{2}$ is the free-particle Dirac operator and $\hat{V}=\hat{J}-\hat{K}+\hat{V}_{\mathrm{nuc}}$ is the SCF one-body potential. The matrix $D_{ij}=\braket{\Phi_{i}}{\hat{\mathfrak{D}}}{\Phi_{j}}$ is the projection of the mean-field Hamiltonian $\hat{\mathfrak{D}}$ in the occupied subspace spanned by $N$ orthonormal spinors. These are a set of coupled first-order differential equations. Its solutions are 4-component spinors; each component is a complex function and the set of solutions describes both positive and negative energy states. We rearrange Eq. (10) as:
@@
The nonrelativistic Schrödinger equation, and its mean-field equivalents, can be solved by iterative convolution with the Green’s function for the shifted Laplacian [41]. Following Blackledge and Babajanov [42], we formulate a similar iterative convolution for Eq. (11). First, we introduce the identity on the left-hand side:
$(\hat{h}_{D}-D_{ii})(\hat{h}_{D}+D_{ii})(\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}$ $=\left[\hat{h}_{D}^{2}-D_{ii}^{2}\right](\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}$ (12)
@@
$\hat{h}_{D}^{2}-D_{ii}^{2}=\left[-c^{2}\nabla^{2}+m^{2}c^{4}-D_{ii}^{2}\right]I_{4}.$ (13)
@@
which can be formally inverted to give:[43]
@@
$G_{i}\equiv\left[-\nabla^{2}+\frac{m^{2}c^{4}-D_{ii}^{2}}{c^{2}}\right]^{-1}I_{4}=I_{4}\frac{1}{4\pi}\frac{\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|}$ (16)
@@
since we are seeking bound states, i.e. $D_{ii}<mc^{2}$. As noted by Anderson et al. [40], one can use the properties of the convolution product to rearrange Eq. (15) as:
@@
See also Appendix A for a detailed derivation of these equivalences. As we show in Appendix B, the iterative convolution of Eq. (18a) uses the inverse-distance-cube kernel, which also appears in the context of the relativistic two-electron gauge potential [44]. As shown by us in Ref. [35], application of this kernel is computationally intensive and not numerically robust. The second form, Eq. (18b), is computationally more efficient: we apply the same convolution kernel used in the nonrelativistic case first, smoothing out any sharp features introduced by multiplication with the potential, and then we compute its derivatives [45]. The final form in Eq. (18c) requires application of the derivative operator on $V\phi_{i}$, which might have non-smooth features and thus introduce numerical errors, convergence slowdown or excessive memory demands.
@@ 3 Developing of Squared Dirac Operator within Multiwavelets Field
Following Wallmeier and Kutzelnigg [4], the square of the Dirac operator is positive and bounded from below, while obviously sharing eigenfunctions and eigenvalues with the original operator:
$\hat{\mathfrak{D}}^{2}\Phi_{i}$ =$=\left\{c^{2}p^{2}I_{4}+c[\vec{\alpha}\cdot\vec{p},\hat{V}]_{+}+m^{2}c^{4}+mc^{2}[\beta,\hat{V}]_{+}+\hat{V}^{2}I_{4}\right\}\Phi_{i}$ (19)
@@
At convergence, the positive-energy orbital energies of $\hat{\mathfrak{D}}$ are $\epsilon_{i}=-mc^{2}+\sqrt{m^{2}c^{4}+2mc^{2}\omega_{i}}$, where $\omega_{i}$ are the eigenvalues of the shifted-and-scaled SCF equations above. Eq. (20) is close to the nonrelativistic Hartree-Fock (HF) equations, despite the appearance of more complicated potential-like terms. Similarly to Eq. (10) it can be recast as an integral equation and solved by iterative convolution:
$\Phi_{i}=-2mG_{i}\star\left[\mathcal{V}\Phi_{i}+\frac{1}{2mc^{2}}\sum_{j\neq i}(\hat{\mathfrak{D}}^{2})_{ij}\Phi_{j}\right],$ (22)
@@
in the converged occupied subspace and for a complete basis (MWs are to be regarded as a complete basis to within their predefined precision), this relation would be exact.111 The MRA construction only explicitly refines the occupied orbital space and the SCF procedure iteratively block-diagonalizes the matrix representation of the Fock operator. Away from convergence, the off-diagonal blocks will be nonzero: $\begin{pmatrix}\mathbf{O}&\mathbf{A}\\ \mathbf{A}^{\dagger}&\mathbf{V}\end{pmatrix}.$ Forming the square of the Fock operator as the product gives: $\begin{pmatrix}\mathbf{O}^{2}+\mathbf{A}\mathbf{A}^{\dagger}&\mathbf{O}\mathbf{A}+\mathbf{A}\mathbf{V}\\ \mathbf{A}^{\dagger}\mathbf{O}+\mathbf{V}\mathbf{A}^{\dagger}&\mathbf{V}^{2}+\mathbf{A}^{\dagger}\mathbf{A}\end{pmatrix}$ showing that only when $\mathbf{A}=0$ (at convergence) the representation in the occupied basis of $\mathbf{F}^{2}$ will be equal to the product of the basis representation of $\mathbf{F}$ with itself. The exponent $\mu_{i}$ in Eq. (17) is now:

### 27 Feb 2026 v4

# Alpert, Beylkin, Ginez and Vozovoi ae almost everywhere

@@
ABGV
Alpert, Beylkin, Ginez and Vozovoi
ae
almost everywhere
@@
DHF
Dirac–Hartree–Fock
DIIS
Direct Inversion in the Iterative Subspace
@@
IEFPCM
Integral Equation Formalism polarizable continuum model (PCM)
@@
RDM
1-body reduced density matrix
RI
resolution of the identity
RKB
restricted kinetic balance
@@
SAPT
symmetry-adapted perturbation theory
@@
SCRF
self-consistent reaction field
SERS
surface-enhanced raman scattering
SI
Supporting Information
@@
TRS
time-reversal symmetry
WP
Work Package
@@
[3]Christian Tantardini [1]Luca Frediani
1]Hylleraas center, Department of Chemistry, UiT The Arctic University of Norway, PO Box 6050 Langnes, N-9037 Tromsø, Norway 2]Algorithmiq Ltd., Kanavakatu 3C, FI-00160, Helsinki, Finland 3]Center for Integrative Petroleum Research, King Fahd University of Petroleum and Minerals, Dhahran 31261, Kingdom of Saudi Arabia
@@
Jacopo Masotti    Roberto Di Remigio Eikås    christiantantardini@ymail.com    luca.frediani@uit.no [ [ [
@@ Abstract
In this contribution we revive an approach to solving the Dirac equation originally proposed by Kutzelnigg which makes use of the squared Dirac operator $\hat{\mathfrak{D}}^{2}$. This approach holds the promise to avoid the negative energy solution because the negative energy spectrum is now “folded” on the positive energy side and at the same time provides a convex equation, which is amenable to a minimization process and increased precision in the final result.
The $\hat{\mathfrak{D}}^{2}$ yields an equation similar to the non-relativistic one, yet in a four-component framework, where Multiwavelet tools and algorithms developed for the non-relativistic case can be employed with minor modifications. On the other hand, the use of Multiwavelets is here essential to achieve the full potential of the approach.
We implemented and validated this approach for one- and two-electron systems with increasing nuclear charge. Numerical tests were performed to gauge the actual precision of the approach with respect to either analytical reference values when possible or numerical results obtained with the GRASP code otherwise.
@@
multiwavelets, Dirac equation, squared-Dirac operator, Kutzelnigg, variational collapse
@@ 1 Introduction
Relativistic effects are not a niche in molecular electronic structure: for heavy elements they can qualitatively change bonding, spectra, and response properties through scalar-relativistic contraction/expansion and spin–orbit coupling. The four-component Dirac equation is the natural starting point for a first-principles description of electrons whenever high precision is required beyond the nonrelativistic Schrödinger equation.[Dirac_1928, Pyykko_2012]
Despite its conceptual clarity, a direct variational treatment of the Dirac operator is delicate because its spectrum is unbounded from below. In finite basis representations this can manifest as the well-known variational collapse and related pathologies when the discretization inadequately separates positive- and negative-energy states.[Mark1980-zz, Wallmeier_1981, Mark_1982, Schwarz_1982, dyall2012relativistic, dyall2009relativistic, dyall2007relativistic, gomes2010relativistic, dyall2006relativistic, visscher1991kinetic, gianturco1971some, liu2010ideas, stanton1984kinetic, tatewaki2003prolapses]
A practical way to restore a well-posed variational framework is to replace the Dirac operator by a spectrally safer surrogate that preserves the physical eigenfunctions. The squared Dirac operator, $\hat{\mathfrak{D}}^{2}$, is bounded from below and admits a standard Rayleigh–Ritz treatment while retaining the same eigenfunctions only with squared eigenvalues.[Wallmeier_1981, Wallmeier1984-ta] This makes $\hat{\mathfrak{D}}^{2}$ particularly attractive when the numerical method is variational, and when one seeks systematic convergence rather than basis-set empiricism.
In parallel, adaptive multiwavelets bases provide a route to all-electron calculations with guaranteed, tunable precision and without conventional basis-set error. They refine locally where the solution demands it as close to nuclei and in bonding regions as in asymptotic tails, while keeping a compact representation elsewhere.[Harrison_2004, Anderson_2019] Recent work has demonstrated that fully numerical four-component Dirac–Coulomb calculations for molecules are feasible in such adaptive MW bases.[Anderson_2019, tantardini2023future] What is still missing is an equally systematic MW realization of $\hat{\mathfrak{D}}^{2}$, which would combine (i) the variational robustness of the squared formulation with (ii) the precision control and adaptivity of multiresolution analysis.
We first review the one- and many-electron theory of the squared Dirac operator. For the many-electron case we consider the instantaneous Coulomb interaction, leading to the usual direct ($\hat{J}$) and exchange ($\hat{K}$) terms at the mean-field level. We then present numerical results for simple one- and two-electron systems to assess stability and attainable precision, and we compare directly against the standard Dirac formulation to highlight where $\hat{\mathfrak{D}}^{2}$ offers advantages in a multiresolution setting.
@@ 2 Theoretical Background
We use indices $u,w\in\{x,y,z\}$ for Cartesian components. The $2\times 2$ matrices $\sigma_{u}$ appearing as anti-diagonal blocks are the Pauli matrices. The $4\times 4$ matrix $\beta$ is defined as:
@@
The large and small components can be transformed into each other. Practically, the exact relationship is only applicable for the free-particle problem. In atomic and molecular applications, one initializes the small components by applying the kinetic balance condition, i.e. the nonrelativistic limit of the exact large-small relationship, to the large components [sun2011comparison].
@@ 2.1 Adaptive Multiwavelets Basis
MRA is a numerical framework devised to solve integral and differential equations with arbitrary, predefined precision.[Alpert_MRA] Standard quantum chemical methods expand the unknown molecular orbitals into fixed atom-centered bases and recast the solution of the Hartree–Fock equations as a matrix eigenvalue problem to determine the expansion coefficients [JensenBook]. multiresolution analysis (MRA)-based methods instead solve directly for the occupied MOs, through adaptive refinement of the expansion basis. The mathematical framework enabling such adaptive algorithms relies on the construction of a sequence of orthogonal subspaces which is dense in the space of square-integrable functions for a user-defined simulation box.
@@
where $l=0,\ldots,2^{n-1}$ is the translation index, identifying functions in the sub-interval $l/2^{n},(l+1)/2^{n}]$. This construction defines a telescopic series of scaling spaces:
@@
$V_{n}^{k}\oplus W_{n}^{k}=V_{n+1}^{k},\>\>\>W_{n}^{k}\perp V_{n}^{k},$ (9)
Several features of a MW approach[Alpert_1993] concur to efficient algorithms with tunable and controlled precision in quantum chemistry[Harrison_2004, Frediani_2013]. The vanishing moments of the wavelet basis leads to compact (small) representation of functions[Alpert_MRA], disjoint support enables adaptivity which enhances compactness, the non-standard (NS)-form of operators[Beylkin101016jacha200708001] effectively decouples scales at the critical step of the application of convolution operators, the availability of separated representation of operators[Beylkin_ACHA] alleviates the so-called “curse of dimensionality”.
In practice, this allows to control the error of a calculation within a range defined by a precision parameter $\epsilon$ which typically varies from $10^{-4}$ (labeled MW4) to $10^{-8}$ (labeled MW8) in relative terms with respect to the norm of a target function. Numerical efficiency is maximized when the order of the polynomial $k$ is chosen such that $k\simeq-\log_{10}\epsilon+3$.
For a MWnn ($n=4,5,\ldots,8$) calculation, convergence is considered achieved once the norm of the orbital update drops below $\epsilon=10^{-n}$. For a variational approach this implies that the error in the energy should be $10^{-2n}$.
However, as we showed in our previous work on the full Breit Hamiltonian [tantardini2023future], requesting tighter precision for the Dirac equation, demands a higher polynomial order than what provided by default by our MW library, mrcpp[bast_2023_7967323].
The interested reader is referred to the available literature for additional details about MRA and corresponding details about how to achieve the requested precision in its practical realizations for quantum chemistry.[Alpert_1993, Beylkin2005-kg, Frediani_2013, Harrison_2003, Anderson_2019]
A feature of a MW representation and related orbital optimization is that the whole $L^{2}$ space is spanned, with the requested precision being the only limitation. As such this is the ideal approach to revive the original idea by Kutzelnigg and Wallmeier in [Wallmeier_1981], according to which, $\hat{\mathfrak{D}}^{2}$ yields more precise results than the standard $\hat{\mathfrak{D}}$. As they point out, this approach requires a nearly complete basis, to make use of a resolution of the identity (RI) technique. That was prohibitive with atomic orbital (AO) bases at the time. Multiwavelets, on the other hand, guarantee completeness to within any numerical precision available and this problem is therefore solved. It is possible that today’s Gaussian Type Orbital (GTO) bases would perform better, although the RI might be plagued by basis-set overcompleteness. MWs instead, provide completeness by construction.
@@ 2.2 Integral equation formulation for the Dirac operator
where $\hat{h}_{D}=c\vec{\alpha}\cdot\vec{p}+\beta mc^{2}$ is the free-particle Dirac operator and $\hat{V}=\hat{J}-\hat{K}+\hat{V}_{\mathrm{nuc}}$ is the self-consistent field (SCF) one-body potential (implicitly multiplied by the identity matrix $\mathbb{I}_{4}$). The matrix $D_{ij}=\braket{\Phi_{i}|\hat{\mathfrak{D}}|\Phi_{j}}$ is the projection of the mean-field Hamiltonian $\hat{\mathfrak{D}}$ in the occupied subspace spanned by $N$ orthonormal spinors. These are a set of coupled first-order differential equations. Its solutions are 4-component spinors; each component is a complex function and the set of solutions describes both positive and negative energy states. We rearrange Eq. (10) as:
@@
The nonrelativistic Schrödinger equation, and its mean-field equivalent, can be solved by iterative convolution with the Green’s function for the shifted Laplacian [Kalos1962-ok]. Following Blackledge2013-jh, Anderson_2019 formulated a similar iterative convolution for Eq. (11). We revise here the main steps of the derivation. First, we introduce the identity on the left-hand side:
$(\hat{h}_{D}-D_{ii})\underbrace{(\hat{h}_{D}+D_{ii})(\hat{h}_{D}+D_{ii})^{-1}}_{\mathbb{I}_{4}}\Phi_{i}$ $=\left[\hat{h}_{D}^{2}-D_{ii}^{2}\right](\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}$ (12)
@@
$\hat{h}_{D}^{2}-D_{ii}^{2}=\left[-c^{2}\nabla^{2}+m^{2}c^{4}-D_{ii}^{2}\right]\mathbb{I}_{4}.$ (13)
@@
which can be formally inverted to give:[Beylkin2021-sk]
@@
$G_{i}\equiv\left[-\nabla^{2}+\frac{m^{2}c^{4}-D_{ii}^{2}}{c^{2}}\right]^{-1}=\frac{1}{4\pi}\frac{\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|}$ (16)
@@
since we are seeking bound states, i.e. $D_{ii}<mc^{2}$. As noted by Anderson_2019, one can use the properties of the convolution product to rearrange Eq. (15) as:
@@
See also Appendix A for a detailed derivation of these equivalences. As we show in Appendix B, the iterative convolution of Eq. (18a) uses the inverse-distance-cube kernel, which also appears in the context of the relativistic two-electron gauge potential [Saue2011-qg]. As shown by some of us in Ref. [tantardini2023future], application of this kernel is computationally intensive and not numerically robust. The second form, Eq. (18b), is computationally more efficient: we apply the same convolution kernel used in the nonrelativistic case first, smoothing out any sharp features introduced by multiplication with the potential, and then we compute its derivatives [Anderson2019-bx]. The final form in Eq. (18c) requires application of the derivative operator on $V\phi_{i}$, which might have non-smooth features and thus introduce numerical errors, convergence slowdown or excessive memory demands.
@@ 3 Developing of Squared Dirac Operator within Multiwavelets Field
Following Wallmeier_1981, the square of the Dirac operator is positive and bounded from below, while obviously sharing eigenfunctions and eigenvalues with the original operator:
$\hat{\mathfrak{D}}^{2}\Phi_{i}$ $=\left\{c^{2}p^{2}I_{4}+c[\vec{\alpha}\cdot\vec{p},\hat{V}]_{+}+m^{2}c^{4}+mc^{2}[\beta,\hat{V}]_{+}+\hat{V}^{2}I_{4}\right\}\Phi_{i}$ (19)
@@
At convergence, the positive-energy orbital energies of $\hat{\mathfrak{D}}$ are $\epsilon_{i}=-mc^{2}+\sqrt{m^{2}c^{4}+2mc^{2}\omega_{i}}$, where $\omega_{i}$ are the eigenvalues of the shifted-and-scaled SCF equations above. Eq. (20) is close to the nonrelativistic Hartree-Fock (HF) equations, despite the appearance of more complicated potential-like terms. Once again, it can be recast as an integral equation and solved by iterative convolution:
$\Phi_{i}=-2mG_{i}\star\left[\hat{\mathcal{V}}\Phi_{i}+\frac{1}{2mc^{2}}\sum_{j\neq i}(\hat{\mathfrak{D}}^{2})_{ij}\Phi_{j}\right],$ (22)
@@
in the converged occupied subspace and for a complete basis (MWs are to be regarded as a complete basis to within their predefined precision), this relation would be exact. [^1] The MRA construction only explicitly refines the occupied orbital space and the SCF procedure iteratively block-diagonalizes the matrix representation of the Fock operator. Away from convergence, the off-diagonal blocks will be nonzero: $\begin{pmatrix}\mathbf{O}&\mathbf{A}\\ \mathbf{A}^{\dagger}&\mathbf{V}\end{pmatrix}$. Forming the square of the Fock operator as the product gives: $\begin{pmatrix}\mathbf{O}^{2}+\mathbf{A}\mathbf{A}^{\dagger}&\mathbf{O}\mathbf{A}+\mathbf{A}\mathbf{V}\\ \mathbf{A}^{\dagger}\mathbf{O}+\mathbf{V}\mathbf{A}^{\dagger}&\mathbf{V}^{2}+\mathbf{A}^{\dagger}\mathbf{A}\end{pmatrix}$ showing that only when $\mathbf{A}=0$ (at convergence) the representation in the occupied basis of $\mathbf{F}^{2}$ will be equal to the product of the basis representation of $\mathbf{F}$ with itself. The exponent $\mu_{i}$ in Eq. (17) is now: